Find the measure of an angle it its measure is four times the measure of it complement. This is our last H.W. problem and we can't figure it out. Please help.
Answer by stanbon(75887)

Find the measure of an angle ifn its measure is four times the measure of its complement. This is our last H.W. problem and we can't figure it out. Please help.
---------
Let the smaller measure be "x".
Then the larger measure is "4x"
EQUATION:
x+4x = 90
5x = 90
x= 18 degees
4x = 4*18 = 72 degrees
==========
Cheers,
Stan H.